OpenStudy (anonymous):

this is purely for fun. 2y=4y what is y

OpenStudy (shamim):

y=0

OpenStudy (anonymous):

anyone think different?

OpenStudy (skullpatrol):

2*0 = 4*0 0 = 0 True

OpenStudy (anonymous):

It took some of the kids in this class 5 minutes and some didnt get it until it was explained. i just wanted to see if the kids in my class are a little stupid or if its really a tough thing to figure out

OpenStudy (skullpatrol):

If the concept of what an equation means is not explained properly it would be hard for anyone.

OpenStudy (anonymous):

not true if you have taken a algebra class and all of us have.

OpenStudy (anonymous):

you solve for y. theres nothing hard about that. but for some reason no one thinks about zero

OpenStudy (skullpatrol):

If you try to divide by y you get 2 = 4.

OpenStudy (anonymous):

fair enough but thats why you have to use common sense and realize that the only way this could be true is if y=0

OpenStudy (skullpatrol):

Algebraically: 2y=4y 2y-2y=4y-2y 0=2y 0/2=2y/2 0=y This is how I would explain it.

OpenStudy (anonymous):

i mean i do see how it could be hard but yes thats what someone did to solve it

OpenStudy (skullpatrol):

The idea of y belongs to {0} is the solution set is hard to explain properly.

OpenStudy (skullpatrol):

It is the value that makes both y=0 and 2y=4y true statements.

OpenStudy (anonymous):

thank you haha. well im going to close this question. but i had a very good conversation on this

OpenStudy (skullpatrol):

thanks for asking :-)